OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x2 - 5x + 5 = 0

OpenStudy (jdoe0001):

\(\bf --2x^2 - 5x + 5 = 0\\ \textit{quadratic formula}\\ \begin{array}{llll} y=&-2x^2+&-5x+&+5\\ y=&\ ax^2+&\ bx+&c\\ &\uparrow&\uparrow&\uparrow\\ &a&b&c \end{array}\qquad x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (jdoe0001):

hmm got an extra minus =) \(\bf -2x^2 - 5x + 5 = 0\\ \textit{quadratic formula}\\ \begin{array}{llll} y=&-2x^2+&-5x+&+5\\ y=&\ ax^2+&\ bx+&c\\ &\uparrow&\uparrow&\uparrow\\ &a&b&c \end{array}\qquad x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (anonymous):

I got \[- \frac{ 5 }{ 2 } \pm \frac{ \sqrt{65} }{ 2 }\]